OpenStudy (anonymous):

What is the vertex of y=x^2+2x-1

OpenStudy (anonymous):

x-coordinate of vertex = -b/2a b = 2, a = 1 -b/2a = -2/2 = -1 Plug in x = -1 into equation y = (-1)^2 + 2(-1)-1 = 1 - 2 -1 = -2 Vertex is (-1, -2)